OpenStudy (anonymous):

which of the following demonstrates periodic behavior? Tennis ball thrown up into the air Movement of cd in cd player Flight of football while tossed Take off and landing of jet plane

OpenStudy (anonymous):

@LastDayWork

OpenStudy (anonymous):

I think its the movement of cd in cd player?

OpenStudy (lastdaywork):

"..think.." o.O Can you describe "periodic behavior" ?

OpenStudy (anonymous):

its like the cycles for a sinusoidal function, like their specific descriptions

OpenStudy (anonymous):

I think its B not sure though

OpenStudy (lastdaywork):

B is correct :) Here's something I found on "Encyclopedia Britannica" "periodic motion, in physics, motion repeated in equal intervals of time."

OpenStudy (anonymous):

cool thank you!

OpenStudy (anonymous):

I have a couple of other questions, would it be better on this same post or different ones?

OpenStudy (lastdaywork):

Your call..I am not much into medals.. ;)

OpenStudy (anonymous):

haha ok on here then :)

OpenStudy (anonymous):

y = csc x does not have zeros. True or False? True? :)

OpenStudy (lastdaywork):

"...does not have zeros" Can you describe the term "zero" ??

OpenStudy (anonymous):

zeros would be where the graph touches x axis

OpenStudy (anonymous):

and the graph of y=csc x does not touch the x axis at any moment

OpenStudy (lastdaywork):

Hence, true :)

OpenStudy (anonymous):

awesome :)

OpenStudy (anonymous):

which of the following shows a change in amplitude (amplitude would be the A in the equation for sinusoid function -- y = Asin(B(x +/- c)) +/- D y = tan2x y = -2-cos(x-pi) y = -2cos6x y= 1+ sin x Therefore, based on the general equation I believe the answer would be C?

OpenStudy (lastdaywork):

Do you know how to differentiate an equation ?

OpenStudy (anonymous):

?

OpenStudy (lastdaywork):

I'll take that as a NO :P "which of the following shows a change in amplitude (amplitude would be the A in the equation for sinusoid function -- y = Asin(B(x +/- c)) +/- D" What do you mean by "change in amplitude" ??

OpenStudy (anonymous):

like the amplitude has changed from the parent graph you know the original equation for tan or sin or cos etc.. like y = cos x ---> y = -2cos6x. The graph has now had a vertical stretch of 2 and been flipped across axis due to the negative sign

OpenStudy (anonymous):

Change of 2 in amplitude that's kinda what I mean. Not really sure how to explain it otherwise.

OpenStudy (anonymous):

the amplitude has changed into a different number or something like that not really sure how to explain it

OpenStudy (lastdaywork):

Okay I got it. Lemme read the question again :)

OpenStudy (lastdaywork):

Yep C is correct.

OpenStudy (anonymous):

 which function does the graph show? y = tan 2x y = 3cos4x y = sin(x+ pi) y = -2-cos(x-pi)